OpenStudy (anonymous):

Just wondering. Is the surface area of a sphere the circumference of the sphere squared?

OpenStudy (anonymous):

I meant the circumference of the regular circle right through the middle. \[2 \pi r\]

OpenStudy (pfenn1):

I made an error in the explanation. It should have been: No, because the circumference is given by\[2\pi r\]and the surface area is given by\[4πr^2\]The circumference squared would be \[4\pi^2 r^2\]

OpenStudy (anonymous):

Area is measured in square units. Volume is measured in cubic units. Surface Area- square units.

OpenStudy (anonymous):

How do you derive the surface area of a circle?

OpenStudy (zzr0ck3r):

calculus

OpenStudy (zzr0ck3r):

you can do it with a double or tripple integral

OpenStudy (zzr0ck3r):

and Im sure many other ways